OpenStudy (anonymous):

use the fundamental identities to simplify the expression. 1/tan^2x+1

OpenStudy (goformit100):

@mathavraj

OpenStudy (anonymous):

\[\sin^2(x)+\cos^2(x)=1\] is the most fundamental one can you see how to get a \(\tan^2(x)\) out of this?

OpenStudy (anonymous):

sec^2x?

OpenStudy (lgbasallote):

^close...if the question is \[\frac{1}{\tan^2 x + 1}\] \(\sec^2 x\) is just the \(\tan^2 x + 1\) part

OpenStudy (anonymous):

hmm, which fundamental identity should i use then? O:

OpenStudy (anonymous):

so it snot necessary to change it"?

OpenStudy (anonymous):

what i meant was, start with \[\sin^2(x)+\cos^2(x)=1\] and to get tangent out of it, recall that \(\tan(x)=\frac{\sin(x)}{\cos(x)}\) so divide everything by \(\cos(x)\) to get the new identity \[\tan^2(x)+1=\frac{1}{\cos^2(x)}\]

OpenStudy (anonymous):

then you have the reciprocal of \(\frac{1}{\cos^2(x)}\) which is \(\cos^2(x)\) and your answer

OpenStudy (lgbasallote):

in other words \[\tan^2 x + 1 = \sec^2 x\] substitute that \[\frac{1}{\tan^2 x + 1} \implies \frac{1}{\sec^2 x} \implies \frac{1}{\frac{1}{\cos^2 x}} \implies \cos^2 x\] just depends on how you view life

OpenStudy (anonymous):

thank you :)